# ViM-VQ: Efficient Post-Training Vector Quantization for Visual Mamba

### v1

@@ Abstract
Visual Mamba networks (ViMs) extend the selective space state model (Mamba) to various vision tasks and demonstrate significant potential. Vector quantization (VQ), on the other hand, decomposes network weights into codebooks and assignments, significantly reducing memory usage and computational latency to enable ViMs deployment on edge devices. Although existing VQ methods have achieved extremely low-bit quantization (e.g., 3-bit, 2-bit, and 1-bit) in convolutional neural networks and Transformer-based networks, directly applying these methods to ViMs results in unsatisfactory accuracy. We identify several key challenges: 1) The weights of Mamba-based blocks in ViMs contain numerous outliers, significantly amplifying quantization errors. 2) When applied to ViMs, the latest VQ methods suffer from excessive memory consumption, lengthy calibration procedures, and suboptimal performance in the search for optimal codewords. In this paper, we propose ViM-VQ, an efficient post-training vector quantization method tailored for ViMs. ViM-VQ consists of two innovative components: 1) a fast convex combination optimization algorithm that efficiently updates both the convex combinations and the convex hulls to search for optimal codewords, and 2) an incremental vector quantization strategy that incrementally confirms optimal codewords to mitigate truncation errors. Experimental results demonstrate that ViM-VQ achieves state-of-the-art performance in low-bit quantization across various visual tasks.
@@ 1 Introduction
Recent advancements have sparked growing research interest in state space models (SSMs). [19]. Modern SSMs are particularly adept at capturing long-range dependencies and benefit from parallel training capabilities. Several SSM-based methods, including the structured state space sequence model (S4) [12] and S4D [13], have been introduced to efficiently handle long sequence data across various tasks, owing to their convolutional computations and near-linear complexity. The recent development of Mamba [11] extends the SSM framework by incorporating time-varying parameters and proposing a hardware-aware algorithm for highly efficient training and inference. This has spurred increasing interest in adapting Mamba to vision tasks [44, 25, 17, 41], giving rise to the concept of Visual Mamba networks (ViMs). Among these ViMs, Vision Mamba (Vim) and VMamba have gained particular attention due to their effectiveness and unique architectures.
@@
Model quantization compresses deep neural networks (DNNs), enabling their deployment and inference on resource-constrained devices with limited computational resources. Quantization methods are generally classified into uniform quantization (UQ) [20, 9, 29, 22, 39, 10, 23, 32] and vector quantization (VQ)  [33, 28, 1, 42, 35, 4, 8, 34]. UQ reduces the precision of network weights, while VQ clusters network weights into codebooks and corresponding assignments. Meanwhile, post-training quantization (PTQ) [27, 31] is the most commonly used technique, as it rapidly calibrates the quantization parameters using a small calibration set, without the need for time-consuming fine-tuning. Both UQ and VQ have been widely applied to convolutional neural networks [30] and Transformer-based networks [37], with research on UQ in ViMs already underway [40, 2]. Nevertheless, the feasibility of achieving extreme compression via VQ for ViMs remains unexplored.
@@
Figure 1: Visualization of the weight distribution and quantization error for the Mamba-based blocks in advanced Visual Mamba network. (a) VMamba [25] comprises numerous Mamba-based [11] blocks (e.g., VSS blocks). (b) The weights of these blocks exhibit significant outliers. (c-d) Although conventional vector quantization (VQ) outperforms uniform quantization (UQ) at low-bit precision, substantial quantization errors still severely degrade network performance. (e) Our ViM-VQ enables the quantized network to achieve performance comparable to that of the original network.
@@
We first identify several critical challenges encountered when applying vector quantization (VQ) to ViMs: 1) Outliers in ViMs’ Weights. As shown in Figure 1, the weights of the Mamba-based [11] blocks in ViMs exhibit significant outliers. The presence of significant outliers leads to considerable quantization errors, resulting in unsatisfactory accuracy. 2) Limitations of Existing VQ Methods. As depicted in Figure 2, existing VQ methods for visual networks suffer from substantial calibration steps, significant GPU memory consumption, and suboptimal performance when applied to multi-layered networks such as ViMs. 3) Truncation Error from Assignment Conversion. Figure 3 illustrates that weight average methods suffer from truncation errors introduced during the soft-to-hard assignment conversion, causing degradation of inference performance relative to calibration.
@@
To overcome these challenges, we propose ViM-VQ, an efficient post-training vector quantization method for Visual Mamba networks (ViMs). ViM-VQ introduces a fast convex combination optimization algorithm that efficiently initializes and updates both the convex combinations and the convex hulls, rapidly searching for optimal codewords while significantly reducing computational and memory overhead. Additionally, ViM-VQ adopts an incremental vector quantization strategy to incrementally confirm optimal codewords to mitigate truncation errors. Our major contributions are summarized as follows:
@@
  • Our ViM-VQ efficiently quantizes ViMs and achieves superior performance while significantly reducing both calibration time and GPU memory requirements.
@@
  • Compared to advanced VQ method, experiments demonstrate that our ViM-VQ achieves better performance in low-bit quantization across various visual tasks.
Figure 2: Comparison of different vector quantization (VQ) algorithms. (a) Conventional VQ algorithms [14, 33, 28] assign the optimal codeword to each weight sub-vector based solely on distance, making them unsuitable for the task. (b) Global weighted average VQ algorithms [1, 42] apply soft assignments to all codewords and leverage differentiable weighted averaging to identify optimal codewords based on the objective function. However, these methods require substantial calibration efforts and memory overhead, particularly in multi-layered structures such as Visual Mamba networks (ViMs) [44, 25]. (c) Local weighted average VQ algorithms [4] mitigate issues of global approaches by applying soft assignments to only a few nearest codewords, yet their performance remains suboptimal due to the limited scope. (d) Our ViM-VQ overcomes the above challenges and adopts a fast convex combination optimization algorithm to efficiently initialize and optimize both the convex combinations and the convex hulls to search for optimal codewords. Furthermore, ViM-VQ employs an incremental vector quantization strategy to incrementally confirm optimal codeword and compensate for truncation errors.
@@ 2 Related Work
Mamba [11] is an advanced structured state space model that significantly enhances the capabilities of state space models (SSMs) for processing sequential data. It transforms the parameters of the structured state space model (S4) [12] into learnable functions, introducing a parallel scanning method to improve efficiency. By overcoming the local perception limitations of convolutional neural networks (CNNs) [30] and the quadratic computational complexity of Transformers [37], Visual Mamba networks (ViMs) [24] have found widespread success across a range of visual tasks. Vision Mamba (Vim) [44] represents the first use of Mamba in computer vision, utilizing bidirectional SSMs for global feature modeling and position embeddings to enhance spatial awareness. Subsequently, VMamba [25] introduces a cross-scan module to address direction-sensitive challenges. LocalMamba [17] improves performance by incorporating local inductive biases, while PlainMamba [41] adopts a non-hierarchical structure for better multi-scale integration. Mamba-ND [21] extends the Mamba architecture to multi-modal data by simply altering the sequence order. Despite the impressive performance of ViMs in computer vision tasks, their large size remains a significant obstacle for deployment on edge devices with limited resources.
@@
Model quantization is an effective technique to compress deep neural networks by reducing their model size and improving inference efficiency. Current quantization methods can be broadly classified into two main categories [35]: uniform quantization (UQ), which converts model weight from high to low precision, and vector quantization (VQ), which decomposes model weight into a codebook and corresponding assignments. Most quantization approaches are implemented via post-training quantization (PTQ) [27, 31], which requires only a small calibration dataset. In the UQ domain, GPTQ [10] proposes a layer-wise quantization method that leverages approximate second-order information to quantize weights with minimal accuracy loss. Recent studies have begun exploring UQ methods specifically tailored for ViMs. PTQ4VM [2] introduces per-token static quantization and joint learning of smoothing scale and step size to mitigate outliers. MambaQuant [40] applies the Hadamard transformation to standardize the variance across channels. In the VQ domain, PQF [28] utilizes rate-distortion theory to reorder weights, thereby reducing clustering errors. DKM [1] integrates a differentiable K-Means formulation into the objective function to preserve the network’s accuracy. JLCM [42] builds upon DKM by adjusting the gradient computation for assignments to avoid incorrect updates. VQ4DiT [4] identifies that fixed initial assignments can lead to suboptimal updates, proposing an improved search algorithm to dynamically select more appropriate codewords. Although these methods have shown promising results in CNNs and Transformers, their performance under extremely low-bit quantization settings remains limited in ViMs. To the best of our knowledge, our approach is the first post-training vector quantization solution explicitly designed for ViMs.
@@ 3.1 Mamba
Mamba [11] is a state space model (SSM) inspired by continuous system, mapping a 1-D sequence $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. This system uses $\mathbf{E}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ as the evolution matrix and $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$, $\mathbf{P}\in\mathbb{R}^{1\times\mathtt{N}}$ as the projection matrices. The continuous system works as follows: $h^{\prime}(t)=\mathbf{E}h(t)+\mathbf{B}x(t)$ and $y(t)=\mathbf{P}h(t)$. Mamba introduces a timescale parameter $\mathbf{\Delta}$ to transform $\mathbf{E}$ and $\mathbf{B}$ to discrete matrices $\mathbf{\overline{E}}$ and $\mathbf{\overline{B}}$. This transformation employs the zero-order hold method, defined as:
@@
$\mathbf{y}$ $=\mathbf{x}*\mathbf{\overline{K}}=\mathbf{x}*(\mathbf{P}\mathbf{\overline{B}},\mathbf{P}\mathbf{\overline{E}}\mathbf{\overline{B}},\dots,\mathbf{P}\mathbf{\overline{E}}^{\mathtt{M}-1}\mathbf{\overline{B}}),$ (3)
@@ 3.2 Visual Mamba Networks
Vision Mamba (Vim) [44] and VMamba [25] are the most commonly used Visual Mamba networks (ViMs) [24]. Inspired by ViT [7] and BERT [5], Vim is the first network to extend Mamba to vision tasks. Vim incorporates position embeddings for annotating image sequences and utilizes bidirectional state space models to efficiently compress visual representations. VMamba achieves superior performance compared to Vim. The architectural overview of the tiny-scale VMamba is illustrated in Figure 1(a). The input image is first partitioned into patches by a stem module, yielding an initial 2D feature map. Without additional positional embeddings, VMamba progressively constructs hierarchical features across multiple stages, each comprising a downsampling layer and several Visual State Space (VSS) blocks analogous to Mamba-based blocks.
@@ 3.3 Vector Quantization
Vector quantization (VQ) is a more flexible approach than uniform quantization (UQ), especially suitable for extremely low-bit quantization scenarios. VQ typically employs clustering algorithms such as K-Means [15] to decompose the weight matrix $\mathbf{W}\in\mathbb{R}^{o\times i}$ into codebook and assignments. Specifically, $\mathbf{W}$ is divided into $d$-dimensional sub-vectors $\mathbf{W}=\{w_{o,i/d}\}$, where $o\times i/d$ is the total number of sub-vectors. These sub-vectors are then represented using a codebook $\mathbf{C}\in\mathbb{R}^{k\times d}$, where $c(k)$ refers to the $k$-th codeword of length $d$. The assignments $\mathbf{A}=\{a_{o,i/d}\}$ indicate the index of the codeword that best represents each sub-vector. Finally, the quantized weight $\widehat{\mathbf{W}}$ is reconstructed by replacing each $w_{o,i/d}$ with $c(a_{o,i/d})$:
@@
Assignments $\mathbf{A}$ can be stored using $\frac{o\times i}{d}\times\log_{2}k$ bits, and codebook $\mathbf{C}$ can be stored using $k\times d\times 32$ bits (typically negligible). Therefore, the quantization bit for quantizing 32-bit origin weights can be computed as $\frac{d\times 32}{\log_{2}k}$.
@@ 4.1 Motivation
Quantizing the weights of Visual Mamba networks (ViMs) to extremely low-bit precision significantly reduces memory usage and accelerates inference, making ViMs suitable for deployment on resource-constrained edge devices. However, as illustrated in Figure 1, ViMs’ weights contain numerous outliers, leading to substantial quantization errors when applying conventional uniform quantization (UQ) or vector quantization (VQ), ultimately leading to model performance degradation. The relatively smaller quantization error observed in VQ indicates its superiority over UQ in extremely low-bit quantization scenarios [35, 4], which motivates our exploration of applying VQ to ViMs. As illustrated in Figure 2, existing VQ approaches exhibit the following limitations when applied to ViMs:
@@
Conventional VQ. Deep compression [14], ATB [33], and PQF [28] fail to adequately adjust the distribution of quantized weights to support extremely low-bit quantization. Consequently, these methods have primarily been effective for smaller networks (e.g., ResNet-18 [16]) and show limited scalability for larger models like ViMs.
@@
Global Weighted Average VQ. DKM [1] and JLCM [42] require substantial GPU memory for storing and updating soft assignment ratios and incur significant calibration steps. These limitations render them impractical for multi-layered architectures such as ViMs. For instance, as indicated in Table 5, using a $256\times 4$ codebook for 2-bit quantization significantly increases the number of learnable parameters by $4256\times o\times i/$ 4 per layer, potentially causing out-of-memory issues on resource-constrained devices.
@@
Local Weighted Average VQ. Although VQ4DiT [4] addresses the memory and computational overhead issues present in earlier methods, its performance remains suboptimal due to the limited scope of candidate codewords (see Figure 3. Moreover, initializing soft assignments with uniform values results in a substantial initial discrepancy between the quantized weights and the original weights, causing gradient oscillations during optimization.
@@
Truncation Error from Assignment Conversion. Figure 3 and Table 5 demonstrate that existing weighted average VQ methods neglect the truncation errors introduced during the conversion from soft to hard assignments when determining the optimal codewords. This oversight leads to noticeable inference accuracy degradation compared to calibration results.
Figure 3: Distributions of the largest soft assignment ratios after applying existing weighted average vector quantization (using 4 soft assignments per sub-vector and a codebook of size $256\times 8$). The largest ratios not close to 1.0 demonstrate two issues: (1) a few optimal codewords lie outside current candidate codewords, leading to suboptimal performance, and (2) directly confirming optimal codewords based on largest ratios leads to significant truncation errors.
@@
Therefore, we propose ViM-VQ, a fast post-training vector quantization method specifically tailored for extremely quantizing ViMs, which consists of two innovative components: 1) To overcome the challenges of high GPU memory usage, multiple calibration steps, and suboptimal performance encountered by weighted average VQ methods, we develop a fast convex combination optimization algorithm. This algorithm initializes convex combinations of several nearby candidate codewords and soft assignment ratios. Both the convex combinations and the convex hulls are rapidly optimized, significantly accelerating the search for optimal codewords. 2) To eliminate the truncation error resulting from assignment conversion, we develop an incremental vector quantization strategy, in which soft assignments with high ratios are transitioned from soft to hard to confirm optimal codewords in each calibration iteration. This strategy ensures that parts still undergoing calibration can compensate for truncation errors, preventing these errors from affecting the entire network.
Figure 4: Incremental vector quantization strategy of ViM-VQ. Soft assignments with high ratios (e.g., $\geq$ 0.99) undergo a soft-to-hard conversion to confirm the optimal codewords, while other components that are still being calibrated compensate for truncation errors.

### v2

@@ Abstract
Visual Mamba networks (ViMs) extend the selective state space model (Mamba) to various vision tasks and demonstrate significant potential. As a promising compression technique, vector quantization (VQ) decomposes network weights into codebooks and assignments, significantly reducing memory usage and computational latency, thereby enabling the deployment of ViMs on edge devices. Although existing VQ methods have achieved extremely low-bit quantization (e.g., 3-bit, 2-bit, and 1-bit) in convolutional neural networks and Transformer-based networks, directly applying these methods to ViMs results in unsatisfactory accuracy. We identify several key challenges: 1) The weights of Mamba-based blocks in ViMs contain numerous outliers, significantly amplifying quantization errors. 2) When applied to ViMs, the latest VQ methods suffer from excessive memory consumption, lengthy calibration procedures, and suboptimal performance in the search for optimal codewords. In this paper, we propose ViM-VQ, an efficient post-training vector quantization method tailored for ViMs. ViM-VQ consists of two innovative components: 1) a fast convex combination optimization algorithm that efficiently updates both the convex combinations and the convex hulls to search for optimal codewords, and 2) an incremental vector quantization strategy that incrementally confirms optimal codewords to mitigate truncation errors. Experimental results demonstrate that ViM-VQ achieves state-of-the-art performance in low-bit quantization across various visual tasks.
@@ 1 Introduction
Recent advancements have sparked growing research interest in state space models (SSMs) [19]. Modern SSMs are particularly adept at capturing long-range dependencies and benefit from parallel training capabilities. Several SSM-based methods, including the structured state space sequence model (S4) [12] and S4D [13], have been introduced to efficiently handle long sequence data across various tasks, owing to their convolutional computations and near-linear complexity. The recent development of Mamba [11] further extends the SSM framework by incorporating time-varying parameters and proposing a hardware-aware algorithm for highly efficient training and inference. This has sparked increasing interest in adapting Mamba to vision tasks [44, 25, 17, 41], giving rise to the concept of Visual Mamba networks (ViMs). Among these ViMs, Vision Mamba (Vim) and VMamba have gained particular attention due to their effectiveness and unique architectures.
@@
Model quantization compresses deep neural networks (DNNs), enabling their deployment and inference on resource-constrained devices. Quantization methods are generally classified into uniform quantization (UQ) [20, 9, 29, 22, 39, 10, 23, 32] and vector quantization (VQ)  [33, 28, 1, 42, 35, 4, 8, 34]. UQ reduces the precision of network weights, while VQ clusters network weights into codebooks and corresponding assignments. Among various quantization paradigms, post-training quantization (PTQ) [27, 31] is a particularly popular technique, as it rapidly calibrates quantization parameters using a small dataset without the need for full retraining or time-consuming fine-tuning. Both UQ and VQ have been widely applied to convolutional neural networks [30] and Transformer-based networks [37], with research on UQ in ViMs already underway [40, 2]. Nevertheless, the feasibility of achieving extreme compression via VQ for ViMs remains unexplored.
@@
Figure 1: Visualization of the weight distribution and quantization error for the Mamba-based blocks within Visual Mamba networks (ViMs). (a) VMamba [25] comprises numerous Mamba-based [11] blocks (e.g., VSS blocks). (b) The weights of these blocks exhibit significant outliers. (c-d) Although conventional vector quantization (VQ) outperforms uniform quantization (UQ) at low-bit precision, substantial quantization errors still severely degrade network performance. (e) Our ViM-VQ restores performance to a level nearly identical to the full-precision model.
@@
We first identify several critical challenges encountered when applying VQ to ViMs: 1) Outliers in ViMs’ Weights. As shown in Figure 1, the weights of the Mamba-based [11] blocks in ViMs exhibit significant outliers. The presence of outliers leads to considerable quantization errors, resulting in unsatisfactory accuracy. 2) Limitations of Existing VQ Methods. As depicted in Figure 2, existing VQ methods for visual networks suffer from substantial calibration steps, significant GPU memory consumption, and suboptimal performance when applied to multi-layered networks such as ViMs. 3) Truncation Error from Assignment Conversion. Figure 3 illustrates that weighted average methods suffer from truncation errors introduced during the soft-to-hard assignment conversion, causing degradation of inference performance relative to calibration.
@@
To address these challenges in quantizing ViMs, we propose ViM-VQ, an efficient post-training vector quantization method for Visual Mamba networks (ViMs). ViM-VQ introduces a fast convex combination optimization algorithm that efficiently initializes and updates both the convex combinations and hulls, rapidly searching for optimal codewords while significantly reducing computational and memory overhead. Additionally, ViM-VQ adopts an incremental vector quantization strategy that progressively confirms optimal codewords, thereby mitigating truncation errors. Our major contributions are summarized as follows:
@@
  • ViM-VQ efficiently quantizes ViMs, achieving superior performance while significantly reducing both calibration time and GPU memory requirements.
@@
  • Experiments demonstrate that, compared to other advanced VQ methods, our ViM-VQ achieves superior performance in low-bit quantization across various tasks.
@@ 2 Related Work
Mamba [11] is an advanced structured state space model that significantly enhances the capabilities of state space models (SSMs) for processing sequential data. It transforms the parameters of the structured state space model (S4) [12] into learnable functions, introducing a parallel scanning method to improve efficiency. By overcoming the local perception limitations of convolutional neural networks (CNNs) [30] and the quadratic computational complexity of Transformers [37], Visual Mamba networks (ViMs) [24] have found widespread success across a range of visual tasks. Vision Mamba (Vim) [44] represents the first use of Mamba in computer vision, utilizing bidirectional SSMs for global feature modeling and position embeddings to enhance spatial awareness. Subsequently, VMamba [25] introduces a cross-scan module to address direction-sensitive challenges. LocalMamba [17] improves performance by incorporating local inductive biases, while PlainMamba [41] adopts a non-hierarchical structure for better multi-scale integration. Mamba-ND [21] extends the Mamba architecture to multi-dimensional data by altering the sequence order. Despite the impressive performance of ViMs in computer vision tasks, their large size remains a significant obstacle for deployment on edge devices with limited resources.
@@
Model quantization is an effective technique to compress deep neural networks by reducing their model size and improving inference efficiency. Current quantization methods can be broadly classified into two main categories [35]: uniform quantization (UQ), which converts model weight from high to low precision, and vector quantization (VQ), which decomposes weight into a codebook and corresponding assignments. Most quantization approaches are implemented via post-training quantization (PTQ) [27, 31], which requires only a small calibration dataset. In the UQ domain, GPTQ [10] proposes a layer-wise quantization method that leverages approximate second-order information to quantize weights with minimal accuracy loss. Recent studies have begun exploring UQ methods specifically tailored for ViMs. PTQ4VM [2] introduces per-token static quantization and joint learning of smoothing scale and step size to mitigate outliers. MambaQuant [40] applies the Hadamard transformation to standardize the variance across channels. In the VQ domain, PQF [28] utilizes rate-distortion theory to reorder weights, thereby reducing clustering errors. DKM [1] integrates a differentiable K-Means formulation into the objective function to preserve the network’s accuracy. JLCM [42] builds upon DKM by adjusting the gradient computation for assignments to avoid incorrect updates. VQ4DiT [4] identifies that fixed initial assignments can lead to suboptimal updates, proposing an improved search algorithm to dynamically select more appropriate codewords. Although these methods have shown promising results in CNNs and Transformers, their performance under extremely low-bit quantization settings remains limited in ViMs. To the best of our knowledge, our approach is the first post-training vector quantization solution explicitly designed for ViMs.
@@ 3.1 Mamba
Mamba [11] is a state space model (SSM) inspired by continuous systems, mapping a 1-D sequence $x(t)\in\mathbb{R}$ to $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. This system uses $\mathbf{E}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ as the evolution matrix and $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$, $\mathbf{P}\in\mathbb{R}^{1\times\mathtt{N}}$ as the projection matrices. The continuous systems work as follows: $h^{\prime}(t)=\mathbf{E}h(t)+\mathbf{B}x(t)$ and $y(t)=\mathbf{P}h(t)$. Mamba introduces a timescale parameter $\mathbf{\Delta}$ to transform $\mathbf{E}$ and $\mathbf{B}$ to discrete matrices $\mathbf{\overline{E}}$ and $\mathbf{\overline{B}}$. This transformation employs the zero-order hold method, defined as:
@@
$\mathbf{y}$ $=\mathbf{x}*\mathbf{\overline{K}}=\mathbf{x}*(\mathbf{P}\mathbf{\overline{B}},\mathbf{P}\mathbf{\overline{E}}\mathbf{\overline{B}},\dots,\mathbf{P}\mathbf{\overline{E}}^{\mathtt{M}-1}\mathbf{\overline{B}}),$
@@ 3.2 Visual Mamba Networks
Vision Mamba (Vim) [44] and VMamba [25] are the most commonly used Visual Mamba networks (ViMs) [24]. Inspired by ViT [7] and BERT [5], Vim is the first network to extend Mamba to vision tasks. Vim incorporates position embeddings for annotating image sequences and utilizes bidirectional state space models to efficiently compress visual representations. VMamba achieves superior performance compared to Vim. It processes input images by first partitioning them into patches through a stem module, yielding an initial 2D feature map. Without additional positional embeddings, VMamba progressively constructs hierarchical features across multiple stages, each comprising a downsampling layer and several Visual State Space (VSS) blocks analogous to Mamba-based blocks.
@@ 3.3 Vector Quantization
Vector quantization (VQ) offers more flexibility than uniform quantization (UQ), making it especially suitable for extremely low-bit quantization scenarios. VQ typically employs clustering algorithms such as K-Means [15] to decompose the weight matrix $\mathbf{W}\in\mathbb{R}^{o\times i}$ into a codebook and assignments. Specifically, $\mathbf{W}$ is divided into $d$-dimensional sub-vectors $\mathbf{W}=\{w_{o,i/d}\}$, where $o\times i/d$ is the total number of sub-vectors. These sub-vectors are then represented using a codebook $\mathbf{C}\in\mathbb{R}^{k\times d}$, where $c(k)$ denotes the $k$-th codeword of length $d$. The assignments $\mathbf{A}=\{a_{o,i/d}\}$ indicate the index of the codeword that best represents each sub-vector. Finally, the quantized weight $\widehat{\mathbf{W}}$ is reconstructed by replacing each $w_{o,i/d}$ with $c(a_{o,i/d})$:
@@
Assignments $\mathbf{A}$ can be stored using $\frac{o\times i}{d}\times\log_{2}k$ bits, and codebook $\mathbf{C}$ can be stored using $k\times d\times 32$ bits (typically negligible). Therefore, the quantization bit for quantizing a 32-bit original weight can be computed as $\frac{\log_{2}k}{d}$.
Figure 2: Comparison of different vector quantization (VQ) algorithms. (a) Conventional VQ algorithms [14, 33, 28] rely on simple distance-based assignment, making them unsuitable for the task. (b) Global weighted average VQ algorithms [1, 42] use and update soft assignments to all codewords and select the maximum as the final assignment. However, they suffer from high calibration and memory overhead, especially for multi-layered structures such as Visual Mamba networks (ViMs) [44, 25]. (c) Local weighted average VQ algorithms [4] reduce overhead by limiting soft assignments to a few nearest codewords but compromise performance. (d) Our ViM-VQ overcomes these limitations by using a fast convex combination optimization algorithm and an incremental vector quantization strategy to efficiently identify optimal codewords while compensating for truncation errors.
@@ 4.1 Motivation
Quantizing the weights of Visual Mamba networks (ViMs) to extremely low-bit precision significantly reduces memory usage and accelerates inference, making ViMs suitable for deployment on resource-constrained edge devices. However, as illustrated in Figure 1, the weights of ViMs contain numerous outliers, leading to substantial quantization errors when applying conventional uniform quantization (UQ) or vector quantization (VQ), ultimately resulting in model performance degradation. The observation that VQ incurs a relatively smaller quantization error underscores its superiority over UQ in extremely low-bit scenarios [35, 4], thus motivating our focus on applying VQ to ViMs. As illustrated in Figure 2, existing VQ approaches exhibit the following limitations when applied to ViMs:
@@
Conventional VQ. Deep compression [14], ATB [33], and PQF [28] fail to adequately adjust the distribution of quantized weights to support extremely low-bit quantization. Consequently, these methods have primarily been effective for smaller networks (e.g., ResNet-18 [16]) and have shown limited scalability for larger models like ViMs.
@@
Global Weighted Average VQ. DKM [1] and JLCM [42] demand not only substantial GPU memory for storing and updating soft assignment ratios but also significant calibration steps. These limitations render them impractical for multi-layered architectures such as ViMs. For instance, as indicated in Table 5, using a 256×4256$\times$ 4256 × 4 codebook for 2-bit quantization significantly increases the number of learnable parameters by $256\times o\times i/4$ per layer, potentially causing out-of-memory issues on resource-constrained devices.
@@
Local Weighted Average VQ. Although VQ4DiT [4] addresses the memory and computational overhead issues present in earlier methods, its performance remains suboptimal due to the limited scope of candidate codewords (see Figure 3). Moreover, their approach of initializing soft assignments with uniform values leads to a substantial initial discrepancy between the quantized and original weights, which can cause gradient oscillations during optimization.
@@
Truncation Error from Assignment Conversion. As demonstrated in Figure 3 and Table 5, existing weighted average VQ methods neglect the truncation error that is introduced when converting from soft to hard assignments to finalize the optimal codewords. This oversight leads to noticeable inference accuracy degradation compared to calibration results.
Figure 3: Distribution of maximum soft assignment ratios in weighted average vector quantization (4 candidates per sub-vector, 256×8256$\times$ 8256 × 8 codebook). When these maximum ratios deviate significantly from 1.0, it indicates two issues: (1) optimal codewords lie outside current candidates, limiting performance; (2) hard assignment conversion causes significant truncation errors.
@@
Therefore, we propose ViM-VQ, a fast post-training vector quantization method specifically tailored for the extremely low-bit quantization of ViMs. Our method consists of two innovative components: 1) To overcome the challenges of high GPU memory usage, extensive calibration steps, and suboptimal performance encountered by weighted average VQ methods, we develop a fast convex combination optimization algorithm. This algorithm operates on convex combinations formed by several nearby candidate codewords and their corresponding soft assignment ratios. Both the convex combinations and the convex hulls are rapidly optimized, significantly accelerating the search for optimal codewords. 2) To eliminate the truncation error resulting from assignment conversion, we develop an incremental vector quantization strategy, in which soft assignments with high ratios are transitioned from soft to hard to confirm optimal codewords in each calibration iteration. This strategy allows the sub-vectors still undergoing calibration to compensate for truncation errors, thereby preventing the accumulation of errors throughout the network.